# Radiation forces and torques in optics and acoustics

### v1

@@ V.4.1 Optics
The optical torque (26) due to the absorption of light carrying spin AM by an isotropic particle was measured in [214] for an absorbing particle trapped in the focus of an elliptically polarized Gaussian beam. When the particle is trapped in the center of a vortex beam carrying both spin and OAM, then the torque also acquires the orbital contribution due to the absorption of the OAM [212, 506]. Interesting macroscopic ‘Beth-type’ measurements of the spin-absorption and OAM-absorption torques were performed in [167, 199, 198].
The anisotropy-induced optical torques (29) and (30) were first predicted theoretically and then measured experimentally in historical works [468, 433, 80, 81, 265] (see Section I.1). A breakthrough in the measurement of these torques was made in the work [213], which observed and measured both the aligning and continuous (spin-conversion) types of torque for laser-trapped birefringent microparticles, as shown in Fig. 18(b). Importantly, for small particles, the effective anisotropy, characterized by the anisotropic polarizability tensor , can originate not only from the material birefringence but also from the anisotropic shape of the particle. Various anisotropic-shape particles, such as micro- and nano-rods, trapped and fast-rotated in focused laser beams are currently attracting rapidly growing attention in the context of optomechanics and levitodynamics [264, 8, 445, 309, 311, 310, 241]. Note that in such systems optically-induced torques often combine several mechanisms, including absorption and anisotropy [450, 388].
Optical torques also appear for various complex-shape objects. In any case these originate from the difference between the angular momenta of the incident and scattered light. For example, optical torques were observed for: 2D-chiral-shape particles [262], cylindrical-lens OAM convertor [67], and q-plates converting spin to OAM [253] [see Fig. 9(e)]. An overview of macroscopic torque experiments can be found in [111].
@@ V.4.2 Acoustics
The acoustic spin-absorption torque (26) was described and observed only recently [488, 543]. This is because only recently it was realized that structured sound-wave fields contain nontrivial polarization and spin (10) properties [488, 101, 103]. The presence of such space-varying spin density in two interfering orthogonal plane sound waves was confirmed in [488] by measuring acoustic torque on a resonant-absorbing ‘meta-atom’, Fig. 18(a) (cf. earlier theoretical considerations in [120, 616]).
While it was believed, until recently, that sound waves carry no spin AM, there were numerous experiments detecting the OAM of acoustic vortex beams via torques on absorbing macroscopic objects [559, 170, 24, 511] and small trapped particles [50, 611, 367] (for theory, see [614]).
Anisotropy-induced torques (29) and (30) on anisotropic-shape particles should appear in acoustics in the same way as in optics [513, 202, 342]. Although there are experiments on acoustofluidic torques and rotations of small floating anisotropic particles [479, 478, 74], they still lack clear theoretical analysis based on the acoustic spin.
@@ V.5 Optical chiral-particle sorting
An important example related to chiral particles and forces is the separation of enantiomers, i.e., particles that are identical except for their chirality. Molecular chirality is very important in pharmacology, as different enantiomers usually have significantly different effects [133, 471]. Optical chiral sorting could be a valuable alternative to the current state-of-the-art chemical methods of separation with the advantage of being independent of the chemical structure of the molecule.
@@
Successful chiral manipulation was demonstrated mostly for chiral microparticles larger or comparable to the wavelength [541, 540, 539, 308, 307, 621, 493], and recently for chiral gold nanoparticles [596]. Although chiral microparticles are too big for the dipole approximation which was used in Eq. (IV.2.1), this equation perfectly captures the main features of the observed dynamics.
@@
In particular, the separation in [541] uses two circularly-polarized counter-propagating beams with opposite helicities (i.e., the same directions of spin), Fig. 19(a). The non-chiral radiation-pressure and recoil forces vanish in such field, whereas the chiral radiation-pressure and recoil forces, Eqs. (IV.2.1) and (35), point along the spin density vector . This results in the opposite motions of the right-hand and left-hand enantiomers. Note that the non-chiral gradient forces, caused by the standing-wave intensity gradients, can be neglected in this experiment as the size of the particles is much larger than the period of the standing wave. Therefore, the same approach cannot be used for smaller particles, such as small pharmaceutical molecules, as they would be trapped by the gradient forces in the standing wave maxima or minima.
@@
Next, the experiments [307, 308] produced separation of chiral liquid crystal microspheres using the chiral helicity-gradient force in Eq. (IV.2.1), see Fig. 19(b). Namely, a sinusoidal helicity-density dependence , with was produced by interfering two slightly non-collinear plane waves with orthogonal linear polarizations [124]. Akin to the previous example, the -components of non-chiral optical forces vanish in such field, while the chiral gradient force pushes the right-hand and left-hand enantiomers in the opposite -directions. The helicity-gradient chiral force acting on chiral nanoparticles was also explored in [596].
@@
There is a number of theoretical considerations of transverse chiral-sorting optical forces for particles near interfaces between two media [565, 257, 618]. One of such scenarios has been implemented experimentally for chiral particles floating on the surface of water [493], which essentially employed the interference of the incident and reflected fields.
Further recent suggestions for the optical chiral sorting involve fields near various optical waveguides [239, 366, 349] and a strong effect produced by the combination of a static electric field and an optical standing wave [125]. In spite of the impressive progress in theoretical and experimental studies of chiral radiation forces, implementation of optical chiral sorting of molecules remains an open challenge.
@@ V.6 Acoustofluidic sorting of biological cells
One of the most important applications of acoustic forces and manipulation is the acoustofluidic sorting of biological cells, see reviews [412, 271, 375, 318, 585, 328, 184]. Acoustofluidics is an interdisciplinary research area that combines acoustics with microfluidics. It offers a non-destructive approach for separating biological particles, cells, and organisms.
@@
The acoustofluidic separation relies on the direct flow of a liquid mixture with cells through a microchannel and then the introduction of ultrasound acoustic waves generated by piezoelectric transducers. The main method, shown in Fig. 20, is based on the excitation of a standing acoustic wave across the microchannel [569, 286, 602, 424, 423, 404, 9]. This wave can be either bulk or surface, i.e. evanescent in the third direction orthogonal to the microchannel and wavevectors. In both cases, the -dependence of the acoustic field is described by Eq. (55) with the microchannel width corresponding to half the wavelength, . Such field has a pressure-field node and velocity-field maximum in the middle of the microchannel, as shown Fig. 20(d).
@@
As we discussed in Section V.3.2, the main acoustic radiation force will be in Eq. (22), because the radiation-pressure force vanishes. This force has two contributions from gradients of the pressure-field and velocity-field intensities, weighted by the monopole and dipole polarizabilities of the particle, respectively. Since the pressure and velocity intensity gradients have opposite directions in the field (55), the two gradient-force contributions can also have opposite directions. The total gradient force (22) is given by
@@
where is the acoustic energy density (7). Thus, the direction of this force depends on the relation between the monopole and dipole polarizabilities of the cell. Note that the difference divided by the cell volume is often referred to in the literature as the acoustic contrast factor . Having two types of cells in the microchannel, with their properties corresponding to and , these will be moved and trapped by the opposite-sign gradient forces (56) in the middle and at the edges of the microchannel, respectively (Fig. 20). After this separation, the two types of cells can be delivered to different reservoirs via forking the central and side parts of the microchannel into three channels. Since the polarizabilities depend on many parameters (the cell size, its density and compressibility, the wave frequency, and parameters of the fluid, see Table 4), one can adjust this method to provide efficient sorting to virtually any two types of cells with different physical properties.
@@
There are also other methods of acoustofluidic sorting: e.g., using travelling waves which can produce noticeable recoil forces for Mie particles with [512, 177, 174, 357, 173, 155, 172, 175]. There is also a number of additional effects which must be taken into account and can be employed for manipulation of bio-particles. These are: secondary radiation forces arising from scattering off other particles, bubbles, and surfaces; acoustic streaming; the Stokes drag force related to the motion of particles with respect to the viscous fluid; etc. [7, 363, 256, 412, 375, 271, 184, 62, 585].
@@ V.7 Pulling forces
One of the most intriguing examples of nontrivial radiation forces is the ‘pulling’ or ‘negative’ force, which can pull the particle toward the source of the incident field (wave beams dragging particles against their propagation direction are also called ‘tractor beams’). This topic has recently attracted great attention, both in optics [143, 522, 405, 464, 118, 292, 495, 127, 207, 425, 625, 337, 331, 269] (for reviews, see [187, 183, 332] and acoustics [364, 615, 591, 166, 382, 383, 376, 564].
@@
The most developed mechanism producing pulling radiation forces uses strong forward scattering, such that the backward-directed recoil force (Sec. III.4) would prevail over the forward radiation-pressure force [118, 143, 405, 522, 615, 166, 376]. Such situation can occur both in the Mie resonant regime, where strong forward scattering is produced by the interference of the excited multipoles (see Figs. 8 and 11), and in the geometrical optics regime of large objects. For example, Fig. 21(a) shows optical pulling/pushing forces dependent on the size of Mie particles [118], whereas Fig. 21(b) shows acoustic pulling force on a macroscopic prism-like target [166]. An important restriction of this approach is that it essentially requires two or more (e.g., in Bessel beams) non-collinear plane waves in the incident field to produce the pulling force on a passive object in a homogeneous medium.
@@
There are other methods to generate pulling forces. First, active particles with gain are characterized by negative imaginary part of the polarizability, . This immediately results in negative radiation-pressure forces and torques in Eqs. (22) and (23). Such pulling forces were considered both in optics [384, 566, 221, 11] and acoustics [443, 444, 376]. Second, various temporal modulations of the incident beam (which becomes poly-chromatic in this case) can produce pulling forces [464, 329, 383]. Such forces can be interpreted as gradient forces from the longitudinal intensity gradients (which can move opposite to the main wavevector [464]). Third, various pulling forces can appear for particles near interfaces between two media [292, 360, 436, 425, 564, 563]. In the dipole approximations, the corresponding forces are described by the formalism of Section IV.7.
@@
Other approaches to radiation pulling forces involve complex structured media [331, 625], nontrivial structured fields with opposite-sign projections of the energy flow and canonical momentum density [326], chirality [127, 207], photophoretic forces [495], complex wavefront shaping (Section IV.8) [269], etc.
@@
Rotational counterparts of pulling forces are negative torques which are directed opposite to the spin of the incident field (at least, in projection on a chosen axis) [253, 142, 254, 186, 491, 434]. For an isotropic sphere in a homogeneous medium the negative torque can be achieved only for active particles with . For passive particles, one needs to break their spherical symmetry to generate higher angular momentum in the scattered light, resulting in a negative recoil torque [253, 142, 254]. For example, an optical q-plate transforming incident circularly-polarized wave into a vortex beam with higher angular momentum experiences a corresponding negative recoil torque [253] (see Fig. 9(e)). Another example of negative optical torque for clusters of nanoparticles [254] is shown in Fig. 18(c). Finally, for finite-size particles, torques can appear not only from the spin of the incident field but also from gradients of the radiation forces, which push different sides of the particle differently. Such ‘orbital’ torques can compete with the spin-induced torques and result in negative torques.
@@ VI Conclusions
In this review, we aimed to provide a comprehensive and unifying picture of optical and acoustic radiation forces and torques. Based on numerous earlier studies, we have formulated a universal approach which naturally covers all the main results in optics and acoustics and allows one to easily apply and cross-apply it to a variety of problems. Moreover, we have described the main examples and applied areas where the radiation forces and torques play an important role. We believe that describing a rich variety of optical and acoustic phenomena within such a universal framework is illuminating and can stimulate further insights, as well as exchange of fruitful ideas between different fields.
Of course optical and acoustic forces are involved in a number of areas (e.g., optical tweezers, acoustofluidics, etc.), where understanding individual details is important for application to specific systems. For these further details and peculiarities we refer the readers to numerous reviews (see Section I.3) and original articles dedicated to specific areas and effects. Our goal here was to provide a ‘big picture’ of the main phenomena. We also aimed to make this review pedagogical, with necessary technical Appendices, so that all the main results can be readily traced and derived by newcomers to the field.
Starting with the quite general case of an arbitrary particle in a an arbitrary inhomogeneous (yet monochromatic) wavefield, we have considered a number of important examples of particles (Rayleigh, Mie, isotropic, anisotropic, chiral, etc.) and fields (vortex, standing, evanescent, etc.), as well as a number of specific effects: optical and acoustic trapping, optical chiral sorting, acoustofluidic sorting, pulling forces, etc.
@@
Coherent use of the fundamental dynamical characteristics of monochromatic wave fields: canonical momentum, spin angular momentum, and energy densities, which have universal mathematical forms for different types of waves.
@@
Calculations of the radiation forces and torques via general integrals of the momentum and angular-momentum fluxes. In most important basic cases, the results are expressed via (A) and (B).
@@
As described in Section I, the association of the radiation forces and torques with the momentum and angular momentum of the incident field is justified starting from very early historical studies, because mechanical actions of the wave fields on matter is the most straightforward way to observe and measure the corresponding dynamical properties of the waves.
@@
Importantly, the framework provided in this review greatly simplifies analysis and understanding of many results in the existing literature, and will undoubtedly facilitate future research. Let us provide a few explicit examples.
@@
First, analyzing the longitudinal and azimuthal components of the Poynting vector in a polarized optical Bessel beam with a vortex charge , [406] found that these components can be negative in some areas and interpreted this as a “negative propagation” in the beam, which can lead to negative scattering (i.e., radiation-pressure, in our terminology) forces. However, our formalism clearly shows that the radiation-pressure forces are determined by the canonical momentum density rather than the Poynting vector, and the longitudinal and azimuthal components of the canonical momentum are always positive in such Bessel beams [231]. The true “negative propagation”, i.e., negative phase gradients, leading to negative radiation-pressure forces, can appear in superposition of Bessel beams or different modes [110, 77, 85, 196, 163, 232].
@@
Second, analyzing acoustic torque on isotropic particles in similar sound Bessel beams, [613] found that the longitudinal component of such torque can surprisingly be negative, in contrast to the positive sign of the OAM number . However, it is clear from our formalism that the torque on a small isotropic particle is determined by the spin density, rather than the OAM, and the longitudinal component of the spin density in acoustic Bessel beams is indeed negative in some areas [101]. Thus, this result does not look so surprising anymore.
@@
Third, the vast literature on acoustic and acoustofluidic forces and torques often operates with rather cumbersome acoustic-force expressions involving the material parameters of specific particles and specific field distributions. Expressing acoustic forces via universal parameters, such as the monopole and dipole polarizabilities of the particle, as well as the fundamental dynamical field characteristics, greatly simplifies the main equations and their understanding. For instance, the acoustic gradient force in the acoustofluidic sorting of molecules using a standing ultrasound wave (Section V.6) is usually written in its final form [424, 271, 318] that does not provide physical intuition about the origin of the result and therefore can readily adopt typos [585]. In contrast, presenting the same equation via the particle polarizabilities and energy-density gradients, as in our review, elucidates the physical origin of the phenomenon and equations [375].
@@
Another advantage of the unifying approach presented in this review is that it can be readily extended to other types of systems. For example, recent experiments showed the presence of the wave-induced forces and torques on floating particles in structured surface water waves [562]. Our approach allows one to readily associate these forces and torques with the generic radiation-pressure and gradient forces, as well as spin-induced torques. Moreover, similar wave-induced forces appear for particles interacting with quantum matter waves [555, 622, 246], and even further – for massive objects interacting with gravitational waves.
@@
Naturally, we could not encompass all the radiation force and torque phenomena. The most obvious restriction, assumed throughout this review, is the use of monochromatic field. Radiation forces and torques in polychromatic waves have been studied only for the simplest case of bi-chromatic fields [249, 251, 576, 130, 415, 305, 218, 571, 503], and it has been shown that such fields can have considerable advantages as compared to the monochromatic case. Furthermore, recently there was a great surge of interest in complex space-time structured polychromatic waves [487, 603], which can bring about novel opto-mechanical and acousto-mechanical phenomena.
@@
We hope that this review will contribute to further development of fascinating areas of fundamental and applied science involving the mechanical action of various wave fields.

### v3

@@ V.4.1 Optics
The optical torque (28) originating from the absorption of light carrying spin AM by an isotropic particle was measured in [227] for an absorbing particle trapped in the focus of an elliptically polarized Gaussian beam. When a particle is trapped in the center of a vortex beam carrying both spin and OAM, the torque acquires an additional orbital contribution due to the absorption of the OAM [225, 559]. Furthermore, macroscopic ‘Beth-type’ experiments directly measuring spin-absorption and OAM-absorption torques were reported in [176, 209, 208].
The anisotropy-induced optical torques (31) and (32) were first predicted theoretically and subsequently measured experimentally in early landmark works [516, 472, 83, 84, 280] (see Section I.1). A breakthrough was achieved in [226], which demonstrated and measured both the aligning and continuous (spin-conversion) torques for laser-trapped birefringent microparticles, see Fig. 18(b). Importantly, for small particles, the effective anisotropy, characterized by the anisotropic polarizability tensor , may originate not only from the material birefringence but also from the anisotropic shape of the particle: see microwave experiment [18] with a dipolar antenna. Various anisotropic-shaped particles, such as micro- and nano-rods, trapped and rapidly rotated in focused laser beams, are currently attracting growing attention in optomechanics and levitodynamics [279, 9, 485, 330, 332, 331, 254]. In such systems, optical torques often combine several mechanisms, including absorption and anisotropy [491, 424].
Optical torques also appear for various complex-shaped objects. In every case, these torques originate from the imbalance between the angular momenta of the incident and scattered fields. Examples include 2D chiral-shaped particles [277], cylindrical-lens OAM converters [70], and q-plates converting spin to OAM [267, 389] [see Fig. 9(e)]. A broader overview of macroscopic torque experiments is provided in [116].
@@ V.4.2 Acoustics
The acoustic spin-absorption torque (28) was described and observed only recently [540, 602]. This is because only in recent years was it recognized that structured sound-wave fields can exhibit nontrivial polarization and spin properties, Eq. (12) [540, 106, 87]. The presence of such spatially varying spin density in two interfering orthogonal plane sound waves was experimentally confirmed in [540] by measuring the acoustic torque on a resonant-absorbing ‘meta-atom’, Fig. 18(a) (cf. earlier theoretical considerations in [126, 681]).
While it had long been believed that sound waves carry no spin AM, numerous experiments demonstrated the OAM of acoustic vortex beams by measuring torques on macroscopic absorbing objects [618, 179, 26, 565] and on small trapped particles [53, 676, 399] (see also theory in [679]).
Anisotropy-induced torques (31) and (32) on anisotropic-shaped particles are expected to arise in acoustics in close analogy with optics [567, 213, 371]. Although there have been experiments on acoustofluidic torques and rotations of small floating anisotropic particles [529, 528, 77], these results still lack a consistent theoretical interpretation based on acoustic spin.
@@ V.5 Optical chiral-particle sorting
An important example involving chiral particles and forces is the separation of enantiomers, i.e., particles that are identical except for their handedness. Molecular chirality plays a key role in pharmacology, since different enantiomers often produce markedly different biological effects [140, 519]. Optical chiral sorting could be a valuable alternative to the current state-of-the-art chemical separation methods, with the advantage of being independent of the specific chemical structure of the molecules.
@@
Experimental demonstrations of chiral manipulation have so far been achieved primarily for chiral microparticles with sizes comparable to or larger than the wavelength [596, 595, 594, 329, 328, 687, 546], and more recently for chiral gold nanoparticles [658]. Although such particles are beyond the range of validity of the dipole approximation used in Eq. (IV.2.1), this equation perfectly captures the main features of the observed dynamics.
@@
In particular, the separation in [596] employed two circularly-polarized counter-propagating beams with opposite helicities (i.e., the same spin direction), Fig. 19(a). In this configuration, the non-chiral radiation-pressure and recoil forces vanish, whereas the chiral radiation-pressure and recoil forces, Eqs. (IV.2.1) and (37), act along the spin density vector . This produces opposite motions of the right- and left-handed enantiomers. For , the direction of the chiral force is controlled by . Note that the non-chiral gradient forces, arising from the standing-wave intensity modulation, can be neglected in this case, since the particles are much larger than the standing-wave period. However, this method cannot be applied to much smaller particles, such as small pharmaceutical molecules, because they would be trapped by the gradient forces at the intensity maxima or minima of the standing wave [689].
@@
Next, the experiments [328, 329] demonstrated the separation of chiral liquid-crystal microspheres using the chiral helicity-gradient force in Eq. (IV.2.1), see Fig. 19(b). Specifically, a sinusoidal helicity-density distribution , with was generated by interfering two slightly non-collinear plane waves with orthogonal linear polarizations [130]. As in the previous example, the -components of non-chiral optical forces vanish in this field, while the chiral gradient force, proportional to , drives the right- and left-handed enantiomers in opposite -directions. The helicity-gradient chiral force acting on chiral nanoparticles was also explored in [658].
@@
Several theoretical works have investigated transverse chiral-sorting forces acting on particles near interfaces between different media [625, 271, 683]. One such scenario was realized experimentally in [546] for chiral particles floating on the water surface, which essentially employed the interference between incident and reflected fields.
More recent proposals for optical chiral sorting involve near-field configurations around optical waveguides [252, 398, 379], as well as a strong effect achieved by first orienting the molecules using a static electric field, and then sorting them with an optical standing wave [131, 132]. The latter approach operates entirely within the electric-dipole regime and relies on nonlinear effects, making it fundamentally different from the other methods discussed here (which are based on electric-magnetic dipole coupling ).
Despite significant theoretical and experimental advances in the study of chiral radiation forces, the implementation of optical chiral sorting of molecules remains an open challenge.
@@ V.6 Acoustofluidic sorting of biological cells
One of the most important applications of acoustic forces is acoustofluidic sorting of biological cells, see reviews [450, 286, 410, 339, 646, 350, 193]. Acoustofluidics is an interdisciplinary research area that combines acoustics with microfluidics, providing a non-destructive approach to separating biological particles, cells, and microorganisms.
@@
Acoustofluidic separation is typically realized by driving a liquid mixture of cells through a microchannel while introducing ultrasound waves generated by piezoelectric transducers. The main method, shown in Fig. 20, relies on the excitation of a standing acoustic wave across the microchannel [629, 303, 666, 463, 462, 441, 10]. This standing wave can be either bulk or surface (i.e., evanescent along the third direction, orthogonal to both the microchannel and the wavevectors). In both cases, the -dependence of the acoustic field is described by Eq. (57), with the microchannel width corresponding to half the wavelength, . Such a field exhibits a pressure-field node and a velocity-field maximum at the center of the channel, as shown in Fig. 20(d).
@@
As discussed in Section V.3.2, the main acoustic force in this configuration is the gradient force in Eq. (24), because the radiation-pressure force vanishes. This force has two contributions: from gradients of the pressure-field and velocity-field intensities, weighted by the monopole and dipole polarizabilities of the particle, respectively. Since the pressure and velocity intensity gradients have opposite directions in the standing-wave field (57), the two contributions may also oppose each other. The resulting total force (24) takes the form
where is the acoustic energy density (9). Thus, the force direction is determined by the balance between the monopole and dipole polarizabilities of the cell. The normalized parameter divided by the cell volume is often referred to as the acoustic contrast factor .
@@
If two types of cells are present in the microchannel, one with and the other with , they will experience gradient forces of opposite signs. As a result, one type is trapped near the center of the channel, while the other is pushed toward the edges, see Fig.20. After this separation, the two types of cells can be directed into different reservoirs by branching the central and side parts of the microchannel into three outlets. Since the polarizabilities depend on multiple parameters (cell size, density, compressibility, acoustic frequency, and fluid properties, see Table 5), this method can be tuned to achieve efficient sorting of virtually any two types of cells with distinct physical characteristics.
@@
There are other approaches to acoustofluidic sorting. For example, traveling waves which can produce significant recoil forces for Mie particles with [566, 186, 183, 388, 182, 164, 181, 184]. There are also additional effects which must be taken into account and can be employed for manipulation of bio-particles. These include: secondary radiation forces arising from scattering off neighboring particles, bubbles, and surfaces; acoustic streaming; and the Stokes drag force associated with particle motion relative to the viscous fluid [8, 395, 270, 450, 410, 286, 193, 65, 646].
@@ V.7 Pulling forces
One of the most intriguing examples of nontrivial radiation forces is the ‘pulling’ or ‘negative’ force, which draws a particle toward the source of the incident field. Wave beams that drag particles against their propagation direction are often referred to as ‘tractor beams’. This phenomenon has recently attracted considerable attention, in both optics [151, 577, 442, 510, 124, 311, 548, 134, 218, 464, 694, 363, 354, 284] (for reviews, see [196, 192, 355] and acoustics [396, 680, 653, 175, 418, 419, 411, 623].
@@
The most established mechanism for generating pulling forces relies on strong forward scattering, such that the backward-directed recoil force (Sec. III.4) dominates over the forward radiation-pressure force [124, 151, 442, 577, 680, 175, 411]. This situation can occur in the Mie resonant regime, where forward scattering is enhanced by interference between the excited multipoles (see Figs. 8 and 11), as well as in the geometrical optics regime of large objects. For example, Fig. 21(a) shows optical pulling and pushing forces depending on the size of Mie particles [124], while Fig. 21(b) illustrates an acoustic pulling force acting on a macroscopic prism-like target [175]. An important limitation of this approach is that it essentially requires two or more (e.g., in Bessel beams) non-collinear plane waves in the incident field to produce a pulling force on a passive object in a homogeneous medium.
@@
There are several other methods to generate pulling radiation forces. First, active particles with gain are characterized by a negative imaginary part of the polarizability, . This directly leads to negative radiation-pressure forces and torques in Eqs. (24) and (25). Such pulling forces have been considered in both optics [420, 626, 234, 12] and acoustics [482, 483, 411]. Second, various temporal modulations of the incident beam (which make it poly-chromatic) can also produce pulling forces [510, 352, 419]. These forces can be interpreted as gradient forces arising from longitudinal intensity gradients, which may act opposite to the main wavevector [510]. Third, pulling forces can occur for particles located near interfaces between two media [311, 392, 475, 464, 623, 622]. In the dipole approximations, these forces can be described using the formalism outlined in Section IV.7.
@@
Other approaches to pulling forces involve complex structured media [354, 694], nontrivial structured fields with opposite-sign projections of the energy flow and canonical momentum density [347], chirality [134, 218], photophoretic forces [548], and complex wavefront shaping (Section IV.8) [284].
@@
Rotational counterparts of pulling forces are negative torques, which act opposite to the spin of the incident field (at least when projected on a chosen axis) [267, 150, 268, 195, 544, 473]. For an isotropic sphere in a homogeneous medium, negative torque can be realized only for active particles with . For passive particles, one must break the spherical symmetry to generate higher angular momentum in the scattered field, resulting in a negative recoil torque [267, 150, 268]. For instance, an optical q-plate that transforms an incident circularly-polarized wave into a vortex beam with higher angular momentum experiences a corresponding negative recoil torque [267] (see Fig. 9(e)). Another example of negative optical torque for clusters of nanoparticles [268] is shown in Fig. 18(c). Finally, for finite-size particles, torques can arise not only from the spin of the incident field but also from gradients of the radiation forces, which push different sides of the particle differently. Such ‘orbital’ torques can compete with the spin-induced torques and may result in negative net torques.
V.8 Artificial intelligence advances
Modern optical and acoustic manipulation presents a number of challenges. Recent advances in artificial intelligence and machine learning have proven effective in several areas [159], including: reducing computational costs by predicting particle trajectories and associated forces [351, 120, 644]; enhancing force-field calibration for optical tweezers [29]; Improving tracking of trapped particles [274, 288]; real-time control of optical tweezers [7]; algorithmic smart design of the environment [358, 357, 488]; designing the trapping field [539, 323, 644, 624] or the particle geometry [290]. Machine learning has also enabled several achievements that were previously out of reach. Notable examples include: the inverse design approaches mentioned above, large-scale defect-free assembly [348, 372, 493], and simultaneous manipulation of many-body systems [527, 407].
@@ VI Conclusions
In this review, we have aimed to provide a comprehensive and unifying picture of optical and acoustic radiation forces and torques. Based on numerous earlier studies, we have formulated a universal approach that naturally encompasses the main results in both optics and acoustics, while allowing straightforward application and cross-application to a wide range of problems. We have also highlighted key examples and application areas where radiation forces and torques play a central role. We believe that such a universal framework is illuminating and can stimulate new insights, as well as foster exchange of fruitful ideas across different fields.
Naturally, optical and acoustic forces are central to a number of applied domains (e.g., optical tweezers, acoustofluidics, etc.), where understanding fine details is essential for addressing specific systems. For such details and system-specific phenomena, we refer the reader to a number of dedicated reviews mentioned in Section I.3. Our goal here has been to provide the ‘big picture’ of the main phenomena. We have also aimed to make this review pedagogical, with technical appendices included where necessary, so that newcomers to the field can readily trace and reproduce the main results.
Starting with the general case of an arbitrary particle in an arbitrary inhomogeneous (yet monochromatic) wave field, we have examined a number of important particle types (Rayleigh, Mie, isotropic, anisotropic, chiral, etc.) and field configurations (vortex, standing, evanescent, etc.), as well as a range of specific effects, including optical and acoustic trapping, optical chiral sorting, acoustofluidic sorting, and pulling forces.
@@
Coherent use of the fundamental dynamical characteristics of monochromatic wave fields: canonical momentum, spin angular momentum, and energy densities. These quantities share universal mathematical forms across different types of waves [87].
@@
Calculations of radiation forces and torques via general integrals of the momentum and angular-momentum fluxes. In the most important basic cases, the results are expressed directly in terms (A) and (B).
@@
As described in Section I, the connection between radiation forces / torques and the momentum / angular momentum of the incident field has been recognized since the earliest historical studies. Indeed, the mechanical action of wave fields on matter provides the most direct way to observe and measure the corresponding dynamical properties of waves.
@@
Importantly, the framework provided in this review greatly simplifies the analysis and interpretation of many results in the existing literature, and will undoubtedly facilitate future research. To illustrate this, we next provide a few explicit examples.
@@
First, analyzing the longitudinal and azimuthal components of the Poynting vector in a polarized optical Bessel beam with a vortex charge , [443] found that these components can be negative in some areas and interpreted this as a “negative propagation” within the beam, which could lead to negative scattering (i.e., radiation-pressure, in our terminology) forces. However, our formalism clearly shows that the radiation-pressure forces are determined by the canonical momentum density rather than the Poynting vector. In such Bessel beams, the longitudinal and azimuthal components of the canonical momentum are always positive [244]. True “negative propagation”, i.e., negative phase gradients, leading to negative radiation-pressure forces, can occur only in superposition of Bessel beams or different modes [115, 80, 89, 206, 172, 245].
@@
Second, analyzing acoustic torque on isotropic particles in analogous sound Bessel beams, [678] found that the longitudinal component of the torque can surprisingly be negative, despite the positive sign of the OAM number . Our formalism makes this result more transparent: the torque on a small isotropic particle is determined by the spin density, rather than by the OAM, and the longitudinal component of the spin density in acoustic Bessel beams is indeed negative in certain regions [106]. Thus, this finding is no longer so surprising.
@@
Third, the vast literature on acoustic and acoustofluidic forces and torques often employs cumbersome expressions involving the material parameters of specific particles and specific field distributions. Recasting acoustic forces in terms of universal parameters, such as the monopole and dipole polarizabilities of the particle together with fundamental dynamical field characteristics, greatly simplifies the main equations and their interpretation. For instance, the acoustic gradient force used in acoustofluidic sorting of molecules with a standing ultrasound wave (Section V.6) is typically written in its final form [463, 286, 339], which provides little physical intuition and can readily adopt typographical errors [646]. In contrast, presenting the same force in terms of particle polarizabilities and energy-density gradients, as done in our review, elucidates the physical origin of the effect and clarifies the underlying equations [410].
@@
Another advantage of the unifying approach presented in this review is its straightforward extensibility to other types of systems. For example, recent experiments have demonstrated wave-induced forces and torques acting on floating particles in structured water-surface waves [621, 212], Fig. 22(b). Within our framework, these forces and torques can be naturally identified as generic radiation-pressure and gradient forces, together with spin-induced torques. Moreover, similar wave-induced forces appear for nanoparticles interacting with quantum matter waves [614, 690, 259], Fig. 22(a). On a much larger scale, massive astronomical objects interacting with gravitational waves should also experience wave-induced forces and torques.
Remarkably, the mechanical actions exerted by electron, optical, acoustic, and water waves nicely complement each other across wavelength ranges, and, hence, across the corresponding sizes of manipulated particles [188], see Table 4. These different types of waves enable manipulation of objects spanning from the atomic angstrom scale to the macroscopic meter scale. Furthermore, the water-wave regime extends to colossal wavelengths km in ocean tidal waves, which can exert forces and torques on entire islands of size km [197]. For any kind of wave, the typical magnitude of the radiation force can be roughly estimated as the product of the wave momentum density, the particle’s geometric cross-section (i.e., its squared linear size), and the wave speed, while the typical torque magnitude is given by the same product divided by the wavenumber. These estimations follow from considering the mechanical action of a circularly polarized plane wave incident on a perfectly absorbing particle.
| Wave type | Particle size |
| Electron waves | – |
| Optics | – |
| Acoustics | – |
| Water waves | – |
@@
Naturally, we could not cover all radiation force and torque phenomena. An important limitation, assumed throughout this review, is the use of monochromatic field. Radiation forces and torques in polychromatic waves have so far been studied mainly for the simplest cases of bi-chromatic fields [262, 264, 636, 137, 453, 326, 231, 631, 556, 425] and fields with slow time modulation of the amplitude [510, 352, 419]. It has been shown that such fields can offer considerable advantages and new functionalities compared to the monochromatic case. Recently, there has been a surge of interest in complex space-time structured polychromatic waves [538, 667], which may bring about novel opto-mechanical and acousto-mechanical phenomena. We have also limited our consideration to linear wave scattering. Nonlinear regimes can offer additional phenomena, such as nonlinearity-induced optical torque [597] and bistable trapping effects [301, 685, 417]. Another emergent direction in optical and acoustic manipulation is the study of complex nonlinear and non-Hermitian dynamics of particles, which plays a crucial role in the stability of trapped particles, particularly in many-body systems [364, 496, 362, 381, 492].
@@
We hope that this review will contribute to the further development of fascinating areas of fundamental and applied science involving the mechanical action of diverse wave fields.